OpenStudy (anonymous):

Find the product (3v+2d)(3v-2d)

OpenStudy (anonymous):

=9v^2-6vd+6vd-4d^2 =9v^2-4d^2

OpenStudy (dbzfan836):

6v

OpenStudy (anonymous):

got canceld due to one term with +ve and then with -ve sign..

OpenStudy (dbzfan836):

use the properties 2d-2d=0 3v+3v=6v

OpenStudy (dbzfan836):

can I haz medal now?

OpenStudy (anonymous):

I don't believe that is the correct approach Dbzfan836

OpenStudy (dbzfan836):

why?